# references.bib

### v1

@@ 4 Lower bounds for homogeneous subalgebras
We describe a reduction $\textsc{LBAhalt}\mathrel{\leq^{\textsf{{L}}}_{\text{m}}}\textsc{CMRS}(\textsf{{Homog}})$, proving hardness. Given a LBA $M$ as in Section 2.2 with states $Q=\{q_{0},q_{1},\dots,q_{s},q_{\text{halt}}\}$ and transition function $\delta$, and an input word $w\in\{\texttt{0},\texttt{1}\}^{*}$ of length $n$. Construct the following set of control and padding variables
$\mathcal{C}_{1}\coloneqq Q,\quad\mathcal{C}_{2}\coloneqq\{h_{0},\dots,h_{n+1}\},\quad\mathcal{P}\coloneqq\Set{x_{i,b}}{0\leq i\leq n+1,\ b\in\Gamma=\{\texttt{0},\texttt{1},{\triangleright},{\triangleleft}\}}.$
We call a monomial of the form $qh_{i}x_{0,{\triangleright}}x_{1,b_{1}}\dotsm x_{n+1,{\triangleleft}}$ correctly formed. Using the transition function $\delta$, we build the replacement rules $\mathcal{R}$ of two kinds:
  1. (1)
    For each transition $\delta(q,b)=(q^{\prime},c,{\rm L})$ add the rules $qh_{i}x_{i,b}-q^{\prime}h_{i-1}x_{i,c}$ for all $i=0,\dots,n+1$. For a right movement $\delta(q,b)=(q^{\prime},c,{\rm R})$ similarly add $qh_{i}x_{i,b}-q^{\prime}h_{i+1}x_{i,c}$.
  2. (2)
    Add transitions $q_{\text{halt}}h_{i}x_{i,b}-q_{\text{halt}}h_{i}x_{i,\texttt{0}}$ and $q_{\text{halt}}h_{i}-q_{\text{halt}}h_{1}$ for all $i=0,\dots,n+1$ and $b\in\Gamma$.
These are homogeneous replacement rules controlled with respect to $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$, and rules of the first kind preserve correctly formed monomials. Let $m_{w}\coloneqq q_{0}h_{1}x_{0,{\triangleright}}x_{1,w_{1}}\dotsm x_{n,w_{n}}x_{n+1,{\triangleleft}}$ and $m_{\text{end}}\coloneqq q_{\text{halt}}h_{1}x_{0,\texttt{0}}\dotsm x_{n+1,\texttt{0}}$. Then the logspace-computable reduction function is
$(M,w)\;\mapsto\;(\mathcal{C}_{1}\mathbin{\dot{\cup}}\mathcal{C}_{2}\mathbin{\dot{\cup}}\mathcal{P},\mathcal{R},m_{w},m_{\text{end}}).$
If $M$ halts on $w$ via a sequence of transitions, then applying the same sequence of rules of the first kind to the monomials we have $m_{w}\equiv_{\mathcal{R}}q_{\text{halt}}h_{i}x_{0,b_{0}}\dotsm x_{n+1,b_{n+1}}$ for some $i$ and $b_{0},\dots,b_{n+1}\in\Gamma$. Applying the rules of the second kind we can clean up the final configuration by replacing all symbols by 0 and moving to position $1$, so $m_{w}\equiv_{\mathcal{R}}m_{\text{end}}$.
Conversely, assume that given $(M,w)$ we have $m_{w}\equiv_{\mathcal{R}}m_{\text{end}}$, we need to show that $M$ eventually reaches $q_{\text{halt}}$. Denote the sequence of monomials and applied replacement rules as
$m_{w}\equiv^{\pm R_{1}}m_{1}\equiv^{\pm R_{2}}\cdots\equiv^{\pm R_{l}}m_{l}=m_{\text{end}},\quad R_{1},\dots,R_{l}\in\mathcal{R},$ (2)
and assume that $l$ is as small as possible. Let $e$ be the smallest index such that $m_{e}$ involves $q_{\text{halt}}$. Since $q_{\text{halt}}\nmid m_{w}$ but $q_{\text{halt}}\mid m_{\text{end}}$, we see that $R_{1},\dots,R_{e}$ must be of the first kind and hence $m_{1},\dots,m_{e}$ are correctly formed. Once we argue that none of the replacements is a reverse step $m_{i-1}\equiv^{-R_{i}}m_{i}$, the sequence (2) translates into a valid sequence of configurations and transitions of the LBA $M$.
So assume that the $i$-th replacement is a reverse step, and assume that $i\leq e$ is maximal with this property. The replacement $R_{e}$ must be a forward step by choice of $e$, hence $i<e$ and we have
$m_{i-1}\equiv^{-R_{i}}m_{i}\equiv^{R_{i+1}}m_{i+1}.$
In particular $m_{i}\equiv^{R_{i}}m_{i-1}$, but since the transition function $\delta$ is deterministic (i. e. the transition relation is right-unique), we have $m_{i-1}=m_{i+1}$, contradicting minimality of $l$, the total length (2).∎
The key idea for reducing CMRS to subalgebra membership is the following: If $m\equiv^{R}m^{\prime}$, then $m-m^{\prime}=R\cdot x^{\gamma}\in\mathbb{K}[\mathcal{P}\cup\mathcal{R}]$, and using the control property one can reverse this procedure. The precise statement is the following lemma, similar to [Mayr1982, Lemma 1 & 2].
{lemma}
Given a controlled monomial replacement system $({\underline{x}}=\mathcal{C}_{1}\mathbin{\dot{\cup}}\dots\mathbin{\dot{\cup}}\mathcal{C}_{m}\mathbin{\dot{\cup}}\mathcal{P},\,\mathcal{R})$ and controlled monomials $x^{\alpha},x^{\beta}\in\textsf{{Mon}}({\underline{x}})$ as in the definition of CMRS. The following are equivalent:
  1. (a)
    $x^{\alpha}\equiv_{\mathcal{R}}x^{\beta}$;
  2. (b)
    $x^{\alpha}-x^{\beta}\in\mathbb{Z}[\mathcal{P}\cup\mathcal{R}]\subseteq\mathbb{Z}[{\underline{x}}]$;
  3. (c)
    $x^{\alpha}-x^{\beta}\in\mathbb{K}[\mathcal{P}\cup\mathcal{R}]\subseteq\mathbb{K}[{\underline{x}}]$.
@@
The implications (a)$\Rightarrow$(b)$\Rightarrow$(c) are straightforward. For the implication (c)$\Rightarrow$(a) let $\mathcal{P}=\{x_{1},\dots,x_{k}\}$, =$\mathcal{R}=\{R_{1},\dots,R_{r}\}$ and let $p=p(t_{1},\dots,t_{k},\tilde{t}_{1},\dots,\tilde{t}_{r})$ be a certificate for membership $g\coloneqq x^{\alpha}-x^{\beta}\in\mathbb{K}[\mathcal{P}\cup\mathcal{R}]$, that is, $g=p(x_{1},\dots,x_{k},R_{1},\dots,R_{r})$. Since $x^{\alpha}-x^{\beta},R_{1},\dots,R_{r}$ are controlled polynomials, setting the variables in $\mathcal{C}$ to zero yields the identity
$0=g(x_{1},\dots,x_{k},0,\dots,0)=p(x_{1},\dots,x_{k},0,\dots,0).$
Hence every term in $p$ involves at least one $\tilde{t}_{i}$, so $g$ is a linear combination of multiples of elements from $\mathcal{R}$, in other words $g\in\langle\mathcal{R}\rangle$. This allows us to apply the aforementioned analogous result for ideals [Mayr1982, Lemma 2], which says that $x^{\alpha}\equiv_{\mathcal{R}}x^{\beta}$ disregarding control variables (so $x^{\gamma}\in\textsf{{Mon}}({\underline{x}})$ in Equation 1). On the other hand, since $x^{\alpha}$ contains exactly one variable from each $\mathcal{C}_{i}$ and the same is true for all rules $\mathcal{R}$, no control variable can be used in $x^{\gamma}$ anyway, hence $x^{\alpha}\equiv_{\mathcal{R}}x^{\beta}$ in the controlled sense too. ∎
Combining these two results gives the lower bound on the complexity of $\textsc{AlgMem}_{\mathbb{K}}$.
@@
$({\underline{x}}=\mathcal{C}_{1}\mathbin{\dot{\cup}}\dots\mathbin{\dot{\cup}}\mathcal{P},\mathcal{R},x^{\alpha},x^{\beta})\;\mapsto\;(\{f_{i}\}_{i}=\mathcal{P}\cup\mathcal{R},g=x^{\alpha}-x^{\beta})$
provides reductions $\emph{{CMRS}}\mathrel{\leq^{\textsf{{L}}}_{\text{m}}}\emph{{AlgMem}}_{\mathbb{K}}$ and $\emph{{CMRS}}(\textsf{{Homog}})\mathrel{\leq^{\textsf{{L}}}_{\text{m}}}\emph{{AlgMem}}_{\mathbb{K}}(\textsf{{Homog}})$. In particular, over any field $\emph{{AlgMem}}_{\mathbb{K}}(\textsf{{Homog}})$ is PSPACE-hard, even when restricting the algebra generators to single variables and binomial of degree $\leq 3$.
@@
It remains to prove a matching exponential space complexity lower bound for the general subalgebra membership problem. In the ideal case, the construction for commutative semigroups in [Mayr1982] uses a reduction from exponentially bounded counter machines, and relies crucially on the construction monomials of size $2^{2^{k}}$ using $\mathcal{O}(k)$ replacement rules. In future work, it might be possible to translate this construction into (a variant of) CMRS, and we conjecture that indeed $\textsc{AlgMem}_{\mathbb{K}}$ is EXPSPACE-complete.
The construction of this section has an interesting consequence outside of computational complexity theory.
@@
$\mathcal{C}=\{q_{0},q_{1}\}$$\mathbin{\dot{\cup}}\{h_{0},\dots,h_{n}\},\quad\mathcal{P}=\{x_{1,\texttt{0}},x_{1,\texttt{1}},\dots,x_{n,\texttt{0}},x_{n,\texttt{1}}\},$
@@
$A$ $\coloneqq\mathbb{K}[f_{1},\dots,f_{5n}]\coloneqq\mathbb{K}[\mathcal{P}\cup\mathcal{R}],$
@@ 5 Monomial subalgebras and SAGBI bases
For NP membership one immediately reduces to the case where the polynomial to test $f$ is a monomial due to the $\mathbb{N}^{n}$-grading. In Equation 3 we have $c_{j}\leq\max_{i}\beta_{i}$, so the bit length of $c_{j}$ is bounded by the bit length of $\beta$. Hence non-deterministically guessing $c$ yields a NP-algorithm.
For NP-hardness one can reduce from the NP-complete problem 1in3Sat (Section 2.2). Indeed, given sets $S_{1},\dots,S_{n}\subseteq\{1,\dots,s\}$, then let $\alpha_{1},\dots,\alpha_{s}\in\{0,1\}^{n}$ be the integer vectors with $(\alpha_{i})_{j}=1$ when $i\in S_{j}$. Set $\beta=(1,\dots,1)\in\mathbb{N}^{n}$. Then Equation 3 encodes exactly the subset sum problem, a solution corresponding to $T=\set{i}{c_{i}=1}$. In this construction all monomials $x^{\alpha_{i}},x^{\beta}$ are square-free. ∎
We see that $\textsc{AlgMem}_{K}(\textsf{{Mon}})$ is NP-complete even for polynomials of bounded degree. The same is true if we instead bound the number of variables – if the exponents are encoded in binary.
@@
Encoding the exponents as binary, the unary case $n=1$ is a direct translation of the SubsetSum problem, which is NP-hard. On the other hand, if the monomials are encoded in unary, then a generating-function approach as in [Kane2017] provides a family of circuits in TC0 deciding $\textsc{AlgMem}_{K}(\textsf{{Mon}}(x_{1},\dots,x_{n}))$. Alternatively one can apply (a variant of) Courcelle’s theorem [Elberfeld2012] to obtain TC0-membership. ∎

### v2

@@ 4 Lower bounds on subalgebra membership
If $g=h_{1}f_{1}+\dots+h_{s}f_{s}$ with $h_{i}\in\mathbb{K}[{\underline{x}}]$, then
$tg=h_{1}\cdot(tf_{1})+\dots+h_{s}\cdot(tf_{s}),$
which certifies that $tg$ is in $\mathbb{K}[tf_{1},\dots,tf_{s},{\underline{x}}]$.
Conversely, assume $tg=p(tf_{1},\dots,tf_{s},x_{1},\dots,x_{n})$ with $p\in R\coloneqq\mathbb{K}[u_{1},\dots,u_{s},x_{1},\dots,x_{n}]$. Declare a (non-standard) grading on $R$ and $S=\mathbb{K}[t,x_{1},\dots,x_{n}]$ by giving $\deg t=\deg u_{j}=1$ and $\deg x_{i}=0$, so
$R_{d}\coloneqq\bigoplus_{\begin{subarray}{c}\alpha\in\mathbb{N}^{s}\\ |\alpha|=d\end{subarray}}u_{1}^{\alpha_{1}}\dotsm u_{s}^{\alpha_{s}}\mathbb{K}[x_{1},\dots,x_{n}],\qquad S_{d}\coloneqq t^{d}\mathbb{K}[x_{1},\dots,x_{n}].$
The evaluation homomorphism $R\to S$ substituting $x_{i}\mapsto x_{i}$, $u_{j}\mapsto t_{j}f_{j}$ respects this grading. Since $tg\in S_{1}$ is homogeneous of degree $1$, decomposing $p$ into its graded components $p=\sum_{d}p_{d}$ ($p_{d}\in R_{d}$) reveals that
$p_{d}(tf_{1},\dots,tf_{s},x_{1},\dots,x_{n})=\begin{cases}p(tf_{1},\dots,x_{n})=tg&\text{if }d=1\\ 0&\text{if }d\neq 1.\end{cases}$
Thus by replacing $p$ by $p_{1}$, we may assume that every term in $p$ involves exactly one $u_{j}$. Setting $t=1$, we read
$p(f_{1},\dots,f_{s},x_{1},\dots,x_{n})=g,$
which, by the structure of $p$, is a representation for $g\in\langle f_{1},\dots,f_{s}\rangle$. ∎
{corollary}
Follow the notation from the previous theorem and set $A\coloneqq\mathbb{K}[tf_{1},\dots,tf_{s},{\underline{x}}]$.
  1. (i)
    The representation degree of $g\in\langle f_{1},\dots,f_{s}\rangle$ is one less than the certification degree of $tg\in A$.
  2. (ii)
    The smallest possible number of non-zero terms of all $h_{i}$ in a representation $g=h_{1}f_{1}+\dots+h_{s}f_{s}$ equals the smallest possible number of terms of a certificate for $tg\in A$.
@@
The forward direction in the previous proof gives that the total degree and number of terms for representation of ideal membership are a lower bound for the respective values for the subalgebra (the total degree increases by $1$ by attaching the $u_{j}$ to the $h_{j}$). Conversely, the grading argument shows that if $p$ is a certificate, then so is its graded component $p_{1}$, which has a lower or equal degree and number of terms. Finally, setting $t\mapsto 1$ in $p_{1}$ gives a representation for $g\in\langle f_{1},\dots,f_{s}\rangle$, showing the desired equalities. ∎
With this result we can deduce matching complexity lower bounds for subalgebra membership and its homogeneous variant. We can rephrase Theorem 4.1 as a complexity-theoretic reduction.
@@
$(f_{1},\dots,f_{s};g\in\mathbb{K}[{\underline{x}}])\;\mapsto\;(tf_{1},\dots,tf_{s},x_{1},\dots,x_{n};tg\in\mathbb{K}[t,{\underline{x}}])$
provides a reduction $\emph{{IdealMem}}_{\mathbb{K}}\mathrel{\leq^{\textsf{{L}}}_{\textup{m}}}\emph{{AlgMem}}_{\mathbb{K}}$. More generally, it provides a reduction $\emph{{IdealMem}}_{\mathbb{K}}(\textup{{C}})\mathrel{\leq^{\textsf{{L}}}_{\textup{m}}}\emph{{AlgMem}}_{\mathbb{K}}(\textup{{C}})$ for any class of polynomials C containing single variables and closed under multiplication with variables.
With this we can complete the proof of our main theorem.
Proof of Theorem 1.1.
The complexity-theoretic upper bounds have been established in Theorem 3.1 and Theorem 3.4. The lower bounds now follow from the hardness of $\textsc{IdealMem}_{\mathbb{K}}$ and its homogeneous variant (Theorem 2.12) in conjunction with the reduction in Theorem 4.2. ∎
As mentioned in the beginning of this section, the complexity lower bounds for ideal membership were achieved by studying the word problem for commutative semigroups. Let =$\mathcal{R}=\{(x^{\alpha_{i}},x^{\beta_{i}})\}_{i=1}^{s}$ be a set of pairs of monomials, called replacement rules. Let $\equiv_{\mathcal{R}}$ be the congruence relation on $\textsf{{Mon}}({\underline{x}})$ generated by
$x^{\gamma}x^{\alpha}\equiv_{\mathcal{R}}x^{\gamma}x^{\beta}\qquad(x^{\alpha},x^{\beta})\in\mathcal{R},\ x^{\gamma}\in\textsf{{Mon}}({\underline{x}}).$
Concretely, two monomials $m,m^{\prime}$ are equivalent if one can be turned into the other by repeatedly replacing a factor $x^{\alpha}\mid m$ by the corresponding $x^{\beta}$ or by replacing $x^{\beta}\mid m$ by $x^{\alpha}$.
Word problem for Commutative Semigroups, CSG
Input: Replacement rules =$\mathcal{R}=\{(x^{\alpha_{i}},x^{\beta_{i}})\}_{i=1}^{s}$,
Monomials $m,m^{\prime}\in\textsf{{Mon}}({\underline{x}})$
Output: Is $m\equiv_{\mathcal{R}}m^{\prime}$?
Mayr & Meyer [Mayr1982] prove EXPSPACE-hardness of CSG and provide a reduction $\textsc{CSG}\mathrel{\leq^{\textsf{{L}}}_{\textup{m}}}\textsc{IdealMem}_{\mathbb{K}}(\textsf{Binom})$ via the map
$(\mathcal{R},m,m^{\prime})\;\mapsto\;(x^{\alpha_{1}}-x^{\beta_{1}},\dots,x^{\alpha_{s}}-x^{\beta_{s}};m-m^{\prime}).$
More precisely, the minimal number of terms in a representation for this ideal membership problem equals the minimal number of replacement steps necessary in showing the congruence $m\equiv_{\mathcal{R}}m^{\prime}$. Using the known double-exponential examples from [Mayr1982], we can prove worst-case certification degree lower bounds for subalgebra membership. For an accessible exposition of the double-exponential lower bounds compare [Bayer1988] or the author’s Master’s thesis [kayser2022groebnercomplexity].
Theorem 4.3.
For given $k$ there exist polynomials $f_{1},\dots,f_{s},g\in\mathbb{K}[x_{1},\dots,x_{n}]$ of degree $\leq 6$, $n=\mathcal{O}(k)$, $s=\mathcal{O}(k)$, such that $g\in\mathbb{K}[f_{1},\dots,f_{s}]$, but every certificate $p\in\mathbb{K}[t_{1},\dots,t_{s}]$ has degree at least $2^{2^{n}}$ and at least $2^{2^{n}}$ terms. These $f_{i}$ and $g$ can be chosen as binomials and monomials.
Proof.
The CSG instances from [Mayr1982] lead to ideal membership of $\mathcal{O}(k)$ binomials in $\mathcal{O}(k)$ variables of degree $\leq 5$ (to obtain this exact degree bound, use the slightly improved degree bound from [Bayer1988]) such that every representation has degree and number of terms bounded below by $2^{2^{n}}$. Applying the reduction from Theorem 4.2, we obtain an instance of $\textsc{AlgMem}_{\mathbb{K}}$, whose smallest certificate $p$ has the same number of terms and greater total degree by Section 4. ∎
@@
Following [Bayer1988], this construction can be easily adapted to prove lower bounds of the form $d^{2^{k}}$ at the expense of increasing the algebra generator degree to $d+\mathcal{O}(1)$. This asymptotically matches the double-exponential upper bounds from Section 3. It would be interesting to get more refined lower bounds in terms of other properties of the algebra. For example, what could analogous statements of Theorem 2.14 and Theorem 2.15 be?
While the worst-case examples for CSG are too complicated to recall here, we can give more details in the homogeneous case. Here the certification degree can not blow up, however, we can still produce $\textsc{AlgMem}_{\mathbb{K}}(\textsf{{Homog}})$ instances that require exponentially many terms in their certificates.
@@
The construction is based on the binary-counting LBA from Example 2.7. In [Mayr1997, Theorem 17] Mayr describes how to reduce LBAhalt to commutative semigroups/ideals, proving PSPACE-hardness of $\textsc{IdealMem}_{\mathbb{K}}(\textsf{{Homog}})$. Applying our reduction (Theorem 4.2), we obtain a binary-counting subalgebra. The configuration $(q,i,b_{0},\dots,b_{n+1})$ of the LBA is represented by the monomial $qh_{i}x_{0,{\triangleright}}x_{1,b_{1}}\dotsm x_{n,b_{n}}x_{n+1,{\triangleleft}}$ and the replacement rules/binomial generators mimic the transition function $\delta$.
@@
$\mathcal{T}=\{q_{0},q_{1}\}$$\mathbin{\dot{\cup}}\{h_{0},\dots,h_{n}\},\quad{\underline{x}}=\{x_{1,\texttt{0}},x_{1,\texttt{1}},\dots,x_{n,\texttt{0}},x_{n,\texttt{1}}\},$
@@
$A$ $\coloneqq\mathbb{K}[f_{1},\dots,f_{5n}]\coloneqq\mathbb{K}[\mathcal{T}\cup{\underline{x}}],$
@@ 5 Monomial subalgebras and SAGBI bases
For NP membership one immediately reduces to the case where the polynomial to test $f$ is a monomial due to the $\mathbb{N}^{n}$-grading. In Equation 1 we have $c_{j}\leq\max_{i}\beta_{i}$, so the bit length of $c_{j}$ is bounded by the bit length of $\beta$. Hence, non-deterministically guessing $c$ yields a NP-algorithm.
For NP-hardness, one can reduce from the NP-complete problem 1in3Sat (Section 2.2). Indeed, given sets $S_{1},\dots,S_{n}\subseteq\{1,\dots,s\}$, then let $\alpha_{1},\dots,\alpha_{s}\in\{0,1\}^{n}$ be the integer vectors with $(\alpha_{i})_{j}=1$ when $i\in S_{j}$. Set $\beta=(1,\dots,1)\in\mathbb{N}^{n}$. Then Equation 1 encodes exactly the question for 1in3Sat, a solution corresponding to $T=\set{i}{c_{i}=1}$. In this construction, all monomials $x^{\alpha_{i}},x^{\beta}$ are square-free. ∎
We see that $\textsc{AlgMem}_{\mathbb{K}}(\textsf{{Mon}})$ is NP-complete even for polynomials of degree $\leq n$. The same is true if we instead bound the number of variables – if the exponents are encoded in binary.
@@
$\texttt{maj}(b_{1},\dots,b_{n})=\begin{cases}1&\text{if }\#\set{i}{b_{i}=1}\geq\frac{n}{2}\\ 0&\text{otherwise.}\end{cases}$
For more information on TC0 and circuit complexity classes, see [Vollmer1999, Chapter 4].
@@
Encoding the exponents as binary, the unary case $n=1$ is a direct translation of the UnboundedSubsetSum problem, which is NP-hard.
On the other hand, if the monomials are encoded in unary, then a generating-function approach as in [Kane2017] provides a family of circuits in TC0 deciding $\textsc{AlgMem}_{K}(\textsf{{Mon}}(x_{1},\dots,x_{n}))$. In more detail, Kane describes TC0 circuits deciding the unary vector-valued subset sum problem, which corresponds to solutions of Equation 1 with $c_{i}\in\{0,1\}$. We can reduce the case of arbitrary $c_{i}\in\mathbb{N}$ to the $\{0,1\}$ case by replacing every generator $x^{\alpha_{i}}$ by $x^{\alpha_{i}},x^{2\alpha_{i}},x^{4\alpha_{i}},\dots,x^{2^{\kappa}\alpha_{i}}$, $\kappa=\lfloor\log_{2}|\beta|\rfloor$. Since the numbers are encoded in unary, the input size is proportional to $C=|\beta|+\sum_{i=1}^{s}|\alpha_{i}|$, and this larger generating set has size $\mathcal{O}(C^{2})$.
Alternatively, one can apply a TC0-variant of Courcelle’s theorem to obtain TC0-membership, compare [Elberfeld2012, Theorem 13]. ∎